Find the standard z- score that corresponds to the following:
80% of the distribution is below(to the left of ) this value

A bag of colored blocks contains the follwowing assortment of colors: red(9), blue (24), orange (17), purple (6), green (0), and yellow (24). Construct the probability distribution for x.

Construct a scatterpoint for the (x,y) values below, and answer the following questions.
x 1,2,3,4,5
y -0.5, - 1.0, -1.5, -2.0, -2.5
What would be the slope of this regression line?
Would the correlation between x and y be positive or negative?
How would you interpret these data in terms of linear regression?
How many green elements are required to make this a legitimate probability distribution if there are a total of 50 elements in this sample?
x red blue orange brown green
P(x) 0.14 0.16 0.28 0.24

Answer by stanbon(57967)

Find the standard z- score that corresponds to the following:
80% of the distribution is below(to the left of ) this value
Ans: z = invNorm(0.8) = 0.8416
----------------------------------------------
A bag of colored blocks contains the following assortment of colors: red(9), blue (24), orange (17), purple (6), green (0), and yellow (24). Construct the probability distribution for x.
P(red) = 9/80 ; P(blue) = 24/80 ; P(orange) = 17/80... etc.
---------------------------------------------
Construct a scatterpoint for the (x,y) values below, and answer the following questions.
x 1,2,3,4,5
y -0.5, - 1.0, -1.5, -2.0, -2.5
---
The least-squares regression equation is y = (-1/2)x + 0
-----------------------------------------------------------------------
What would be the slope of this regression line?::: -1/2
-----------------------------------
Would the correlation between x and y be positive or negative? :: negative
------------------------------------
How would you interpret these data in terms of linear regression?
As x increases by 2, y decreases by 1.
===========================================================================
How many green elements are required to make this a legitimate probability distribution if there are a total of 50 elements in this sample?
x red blue orange brown green
P(x) 0.14 0.16 0.28 0.24
---
Let P(green) = 0.18 then # of green = 0.18*50 = 9
=======================================================
Cheers,
Stan H.
==============